A storage battery of EMF 8V, internal resistance 1 ohm is being charged by 120 V D.C. source using a 15 ohm resistor in series in the circuit. Calculate (i)current in the circuit (ii)terminal voltage of the battery (iii)chemical energy stored in the battery in 5 minutes

Grade:12

1 Answers

Vikas TU
7 years ago
Since the battery would in parallel to the 8v and internal resistance 1 ohm and the 15 ohm resistance with series,
the battery purpose would only have to charge the 8 volt Emf, Hence the current drawn in the circuit would be:
I = 8/(1+15) = 0.5 A.
 
Terminal battery Voltage = > 8 – Ir => 8 – 0.5*1 = > 7.5 Voltage.
Heat = I^2(Rnegt)*t => 0.25*16*5*60 => 1200 Joule.